# On the foundations and applications of Lorentz-Finsler Geometry

### v1

@@ Abstract.
The reviewed topics include (i) basics on the setting of cones, Finsler and Lorentz-Finsler metrics and their (nonlinear, anisotropic and linear) connections, (ii) the global structure of Lorentz-Finsler manifolds and its space of null geodesics, (iii) links among Riemannian, Finsler and Lorentz geometries, (iv) real world applications for wildfires and seisms, and discretization in classical and relativistic settings with quantum prospects, and (v) Finslerian variational approach to Einstein equations. The new results include the splitting of globally hyperbolic Finsler spacetimes, in addition to the analysis of several extensions as the case of timelike boundaries.
@@ 1. Introduction
This survey aims to provide an overview of these advances —necessarily non-exhaustive. Readers are assumed to have a basic background in Riemannian Geometry and elementary relativistic concepts, but not necessarily prior training in Finsler Geometry.
@@
The article is structured into five main sections, in addition to this Introduction. Section 2 provides a summary of foundational Finsler concepts and their extension to the Lorentz-Finsler setting. Sections 36 develop four broad topics, which can be read in an essentially independent way. Sections 3 and 4 are more geometrically oriented, covering global Finsler structures and applications to classical geometries. Sections 5 and 6 focus on applications, including real-world phenomena and Finslerian approaches to gravity.
In Section 2, we adopt a pedagogical, non-technical approach to introduce key concepts such as cone structures $\mathcal{C}$, Lorentz-Finsler metrics $L$, and various types of connections, including the anisotropic connection —a notion intermediate between non-linear and Finsler connections. Intuitively, anisotropic connections are the “horizontal part” of Finsler ones and, although not yet mainstream in Finslerian literature, they offer offer a natural way to define covariant derivatives compatible with non-linear connections. At any rate, anisotropic connections turn out to be equivalent to vertically trivial Finsler ones, which may be more familiar to Finsler-oriented readers.
@@
Several choices must be made among the many possibilities found in the literature. Our guiding principle is to establish a reasonable foundational framework that can be adapted for specific studies (e.g., low regularity, singularities, physical models). Some of our choices, further justified throughout the article, include: (i) smooth elements, typically at least $C^{2}$ excluding 0, thus avoiding singular directions (but see for example footnote 24); (ii) connected manifolds with semi-Finsler metrics, that is, requiring the fundamental tensor $g$ is non-degenerate and of constant signature (in parallel with the semi-Riemannian framework of O’Neill [137]); (iii) strong convexity of Finsler indicatrices; (iv) in Lorentzian signature, restriction to causal vectors within a cone structure to focus on causality and distance-maximizing properties, avoiding the arbitrariness of metric extension beyond the cone; (v) retention of the timelike/spacelike terminology in the semi-Finsler case (as in [137]), although this distinction is purely conventional for non-Lorentzian signatures.
@@
Section 3 develops globally hyperbolic Finsler spacetimes. Notice that, as in the Lorentz case, Finsler global hyperbolicity draws some analogies with metric Riemannian completeness. In §3.1, we focus on their global structure and prove the possibility to split it in an orthogonal way, extending the Lorentz case (Theorem. 3.1); this includes the case of spacetimes-with-timelike-boundary (Theorem. 3.3). With this aim, we briefly review the techniques used so far and explain detailedly how those in [24, 25, 3] permit to extend the results to the Lorentz-Finsler case, including additional possibilities of these splittings. In §3.2, following recent paper [84], we explain how the space of null geodesics can also be extended to the Finsler case, thus opening new questions of interest, especiallly when a timelike boundary exists. Finally, Finslerian extensions of Lorentzian singularity theorems are also briefly reviewed.
@@
Section 5 begins with a review of the Fermat principle, Zermelo navigation problem and Snell’s law within the general framework of cone structures, §5.1. This yields multiple applications, among them wildfire and seismic monitoring, §5.2. Furthermore, the discretization of these classical problems opens avenues for applications in Numerical Relativity and other foundational theories in Physics.
@@ 2. Setting on cones, Lorentz-Finsler metrics and connections
The main references for this section are [99, 121, 102].
@@ 2.1. Minkowski, wind Minkowski and Lorentz-Minkowski norms
In this elementary part, we follow mainly [99], which uses Euclidean background, conventions and intuitions in previously posted references [96] and [39]. It is worth pointing out that some low regularity issues on norms are studied in the more general framework of Banach spaces [47] but we will not deepen on this.
@@ 2.1.1. Emergence of non-symmetric norms from anisotropic propagation
Consider a moving object or a propagating wave respect to an observer in a classical (non-relativistic) setting, for example, a zeppeling or sound wave in a (possibly windy) air respect to an observer on earth. The velocity depends on the oriented direction, thus, at each point, the space of (maximum) velocities of propagation yields a closed (compact without boundary) hypersurface $S$, thus, enclosing an open domain $D$. Let us assume that $D$ includes the velocity 0 (in our example, this means that the wind is not too strong, we will discuss this case later, see §2.1.3). Notice that $S$ may be non-symmetric respect to 0, as travelling in the oriented direction of the wind may be faster than on the reversed one. We will be interested the infimum time to arrive from a first point $x$ to a second one $\bar{x}$. When the anisotropic velocities are independent on the point, a reasoning in affine geometry (as around Fig. 4 below), shows that the original $S$ can be effectively replaced by its convex hull (Fig. 1) Consistently, we will assume that $S$ is convex in what follows.
@@
To check that the (strict) triangle inequality is equivalent to the (strict) convexity of $S$, a geometric proof for unit vectors $v,w$ is depicted in Fig. 4 (see also, for example, [96, Prop.2.3 ] or [7] for infinite dimension).
Figure 4. (Strict) convexity implies that the center of the paralelogram with sides $v,w$ lies in (the interior of) $\bar{D}$, thus yielding (strict) triangle inequality.
@@ 2.1.2. Minkowski norms
For the necessary condition, use that the Hessian of $\parallel\cdot\parallel^{2}/2$ at 0, is well defined, symmetric and it agrees with $\parallel\cdot\parallel^{2}$, see [165, Prop. 4.1]. ∎
Then a natural regularity for $\parallel\cdot\parallel$ is smoothness (at least $C^{2}$) away from 0 and, then, so will be its indicatrix $S=\parallel\cdot\parallel^{-1}(1)$. In this case, putting $L_{0}:=\parallel\cdot\parallel^{2}/2$ the fundamental tensor of the norm is:
$g_{v}:=\frac{1}{2}\,\hbox{Hess}\,_{v}L_{0}\qquad\forall v\neq 0.$ (1)
@@
A Minkowski norm is a non-symmetric norm (see Prop. 2.1) which is smooth outside 0 and has a strongly convex indicatrix $S$, that is, its fundamental tensor defined in (1) is positive definite.
The following property of Minkowski norms will rely on our choice of convexity. For any non-symmetric norm smooth away from 0, the Legendre or (non-linear) flat map is
$\begin{array}[]{ccc}\flat:V\rightarrow V^{*},&&v\mapsto v^{\flat}:=(dL_{0})_{v}=g_{v}(v,\cdot).\end{array}$ (2)
@@
This map is clearly continuous, positive homogeneous and onto. It also extends the standard lowering index map of scalar products. However it may be non linear and:
@@ 2.1.3. Wind norms
Coming back to our original problem on propagation §2.1.1, we can wonder what happens if the 0 velocity is not included in the open domain $D$, that is, in the cases of either critical wind, when $0\in S=\partial D$ or strong wind, when $0\not\in\bar{D}:=D\cup\partial D$. In the latter case, a natural cone tangent to $S$ emerges and $S$, up to the tangency points, is the union of two open connected subsets $S_{\hbox{\small{conv}}}$ and $S_{\hbox{\small{conc}}}$ (see Fig. 6).
In a natural way, $S_{\hbox{\small{conv}}}$ can be regarded as the indicatrix of a conic Minkowski norm defined only in the interior of the cone. It also gives a strict triangle inequality for vectors in the interior domain of the cone (see Def. 2.14 and Section 2 in [96]). Analogously, $S_{\hbox{\small{conc}}}$ gives a Lorentz norm and a similar reverse triangle inequality holds [99, §3.1, Appendix B]. In both cases, we can define the square of the norm $L_{0}$ and, thus, the fundamental tensor $g_{v}$ (as in (1)) for $v\neq 0$ in the open conic region, by means of the positive homogeneous extension:
@@
In the Lorentz case, the reverse triangle inequality has a natural interpretation even in a classical setting. Namely, in our example, the Lorentz norm correspond to the case when the zeppelin engine is trying to move in the direction opposite to the wind (describing a straight line in the conic region) but it is drifted by the wind, so that the arrival time is maximized.
@@
Figure 6. Case of strong wind: emergence of convex and concave conic indicatrices in a cone domain.
@@
A wind Minkowski norm on $V$ is a smooth connected compact strongly convex hypersurface $S(=\partial D$). Depending on if $0\not\in\bar{D}$, $0\in S$ or $0\in D$ the wind is called strong, critical or weak, resp. (the latter also corresponding to a Minkowski norm).
@@ 2.1.4. Rheonomic emergence of cones
    Conic: $v\in\mathcal{C}_{0}\Rightarrow\lambda v\in\mathcal{C}_{0},\forall\lambda>0$
@@
    Convex interior: $\mathcal{C}_{0}=\partial A_{0}$ in $V\setminus\{0\}$ with $A_{0}\subset V\setminus\{0\}$ open, convex subset (as around Fig. 3), called the interior of the cone.
@@
  • A vector hyperplane $H\subset V$ is called lightlike (resp. timelike, causal or spacelike) if $H$ is tangent to $\mathcal{C}_{0}$ (resp. contains a timelike vector, lies in one of the previous cases, lies in none of them, resp.)
  • A one-form $\omega:V\rightarrow\mathds{R}$ is called timelike (resp. lightlike, causal) if $\omega(w)>0$ for all $w\in\bar{A}$ (resp. $\omega(w)>0$ for all $w\in A$ and ker $\omega$ is a hyperplane tangent to $\mathcal{C}_{0}$).
@@
(1) The assumptions in the definition of cone can be widely optimized, see [99, §2.1].
@@
(3) Later, causality relations for $\mathcal{C}_{0}$ will behave as if v-v were “past directed”. As a consequence, the natural definition for a vector $v$ (in $V$ or $V\setminus\{0\}$) to be spacelike would be that neither $v$ nor v-v is causal. [^2] Indeed, the word spacelike was not used in the text of [99], but we think it can be clarifying here.
(4) One can check that a vector hyperplane $H$ is spacelike if and only if for one (and then any) causal vector $v$ the affine hyperplane $v+H$ intersects transversely all the directions of $\mathcal{C}_{0}$ (then, yielding a strongly convex hypersurface in $H$ whenever $n>2$); morevoer, any cone must admit spacelike hyperplanes [99, Lemma 2.5, Prop. 2.6].
@@ 2.1.5. Lorentz-Minkowski norms
In §2.1.3, the possibility to define a Lorentz norm associated with a concave indicatrix and a reverse triangle inequality was explained. In the particular case considered therein, the fundamental tensor $g_{v}$ had Lorentzian signature but it was defined only in the (timelike) directions in the interior of the cone (see (3) and Fig. 6). Indeed $L_{0}$ could not be extended continuously by 0 to the cone $\mathcal{C}_{0}$, because the indicatrix was tangent to the cone there. Nevertheless, in the case of a Lorentzian scalar product the unit timelike vectors in one of the cones $\mathcal{C}_{0}$ do yield a Lorentz norm which can be extended by 0 to $\mathcal{C}_{0}$ and such that the corresponding fundamental tensor $g_{v}$ is well defined on $\mathcal{C}_{0}$ and non-degenrate (thus, with Lorentzian signature therein). Such properties are trivial in the case of a scalar product because this product coincides with $g_{v}$ whenever defined, but they serve as a guide for Lorentz norms with convenient regularity.
@@
Again, the hypotheses are not optimized, namely, (3) implies that $\mathcal{C}_{0}$ is a (strongly convex) cone even if one only assumes that $\mathcal{C}$ is the boundary of a domain invariant by positive homotheties (see [99, §3.1]). In any case, the indicatrix $S=L^{-1}_{0}(1)$ will behave qualitatively as in Lorentz-Minkowski and it is asymptotic to the cone $\mathcal{C}_{0}$ (Fig. 8).
@@
Let $L_{0}:\bar{A}_{0}\rightarrow\mathds{R}$ be a Lorentz-Minkowski norm and let $\bar{A}^{\flat}$ be the set of all the causal forms (see Def. 2.6). Its Legendre or flat map is defined formally as in (2) adapting the domain and codomain, that is:
$\flat:\bar{A}_{0}\rightarrow\bar{A}^{\flat}\qquad(dL_{0})_{v}=v\rightarrow v^{\flat}:=g_{v}(v,\cdot).$
Reasoning as around (2), the indicatrix $S$ yields now a positive-homogeneous diffeomorphism between timelike vectors and forms which extend to lightlike ones, making the inverse $\sharp$ well defined.

### v2

@@ Abstract.
The reviewed topics include (i) basics on the setting of cones, Finsler and Lorentz-Finsler metrics and their (nonlinear, anisotropic and linear) connections, (ii) the global structure of Lorentz-Finsler manifolds and its space of null geodesics, (iii) links among Riemannian, Finsler and Lorentz geometries, (iv) applications in classical settings as wildfires and seisms propagation, and discretization in classical and relativistic settings with quantum prospects, and (v) Finslerian variational approach to Einstein equations. The new results include the splitting of globally hyperbolic Finsler spacetimes, in addition to the analysis of several extensions of the Lorentz setting, as the case of timelike boundaries.
@@ 1. Introduction
This survey aims to provide an overview of these advances —necessarily non-exhaustive. Readers are assumed to have a basic background in Riemannian Geometry and elementary Lorentz/relativistic concepts, but not necessarily prior training in Finsler Geometry.
@@
The article is structured into five main sections, in addition to this Introduction. Section 2 provides a summary of foundational Finsler concepts and their extension to the Lorentz-Finsler setting. Sections 36 develop four broad topics, which can be read in an essentially independent way. Sections 3 and 4 are more geometrically oriented, covering global Finsler structures and applications to classical geometries. Sections 5 and 6 focus on applications, including phenomena in applied classical Physics and Finslerian approaches to gravity.
In Section 2, we adopt a pedagogical, non-technical approach to introduce key concepts such as cone structures $\mathcal{C}$, Lorentz-Finsler metrics $L$, various types of connections and the viewpoints of Hamiltonian Mechanics and Contact Geometry. In particular, we stress the viewpoint of anisotropic connections, developed by MA Javaloyes and coworkers [98, 99, 110, 111], which is intermediate between non-linear and Finsler connections and is used in some of the references.
@@
Several choices must be made among the many possibilities found in the literature. Our guiding principle is to establish a reasonable foundational framework that can be adapted for specific studies (e.g., low regularity, singularities, physical models). Some of our choices, further justified throughout the article, include: (i) smooth elements, typically at least $C^{2}$ thus avoiding singular directions (but see for example footnote 24); (ii) connected manifolds with semi-Finsler metrics, that is, requiring the fundamental tensor $g$ is non-degenerate and of constant signature (in parallel with the semi-Riemannian framework of O’Neill [150]); (iii) strong convexity of Finsler indicatrices; (iv) in Lorentzian signature, restriction to causal vectors within a cone structure to focus on causality and distance-maximizing properties, avoiding the arbitrariness of metric extension beyond the cone; (v) retention of the timelike/spacelike terminology in the semi-Finsler case (as in [150]), although this distinction is purely conventional for non-Lorentzian signatures.
@@
Section 3 develops globally hyperbolic Finsler spacetimes. Notice that, as in the Lorentz case, Finsler global hyperbolicity draws some analogies with metric Riemannian completeness. In §3.1, we focus on their global structure and prove the possibility to split it in an orthogonal way, extending the Lorentz case (Theorem. 3.1); this includes the case of spacetimes-with-timelike-boundary (Theorem. 3.3). With this aim, we briefly review the techniques used so far and explain detailedly how those in [26, 27, 4] permit to extend the results to the Lorentz-Finsler case, including additional possibilities of these splittings. In §3.2, we consider the space of null geodesics, introduced in the Lorentz case by R. Low [121] following seminal ideas by R. Penrose [152]. J. Hedicke’s PhD thesis [85] studied it in the Lorentz-Finsler case and the recent progress includes the case with timelike boundary [91]. Finally, Finslerian extensions of Lorentzian singularity theorems are also briefly reviewed.
@@
Section 5 begins with a review on the Fermat principle, Zermelo navigation problem and Snell’s law within the general framework of cone structures, §5.1. This yields multiple applications, among them wildfire and seismic monitoring, §5.2. Furthermore, the discretization of these classical problems opens avenues for applications in Numerical Relativity and other foundational theories in Physics.
@@ 2. Setting on cones, Lorentz-Finsler metrics and connections
The main references along this section are [107, 110, 85, 131] and [52] (the latter considers only the Finsler case, but it is easily adaptable to the Lorentz-Finsler one).
@@ 2.1. Minkowski, wind Minkowski and Lorentz-Minkowski norms
In this elementary part, we follow mainly [107], which uses Euclidean background, conventions and intuitions coming from [104, 41]. It is worth pointing out that some low regularity issues on norms are studied in the more general framework of Banach spaces [53] but we will not deepen on this.
@@ 2.1.1. Emergence of non-symmetric norms from anisotropic propagation
Consider a moving object or a propagating wave respect to an observer in a classical (non-relativistic) setting, for example, a zeppeling or sound wave in a (possibly windy) air respect to an observer on earth. The velocity depends on the oriented direction, thus, at each point, the space of (maximum) velocities of propagation yields a closed (compact without boundary) hypersurface $S$, thus, enclosing an open domain $D$. Let us assume that $D$ includes the velocity 0 (in our example, this means that the wind is not too strong, we will discuss this case later, see §2.1.3). Notice that $S$ may be non-symmetric respect to 0, as travelling in the oriented direction of the wind may be faster than on the reversed one. We will be interested in the infimum time to arrive from a first point $x$ to a second one $\bar{x}$. When the anisotropic velocities are independent on the point, a reasoning in affine geometry (as around Fig. 4 below), shows that the original $S$ can be effectively replaced by its convex hull (Fig. 1). Consistently, we will assume that $S$ is convex in what follows.
@@
To check that the (strict) triangle inequality is equivalent to the (strict) convexity of $S$, a geometric proof for unit vectors $v,w$ is depicted in Fig. 4 (see also, for example, [104, Prop.2.3 ] or [9] for infinite dimension).
Figure 4. (Strict) convexity implies that the center of the paralelogram with sides $v,w$ lies in (the interior of) $\bar{D}$, thus yielding (strict) triangle inequality. analogous picture
@@ 2.1.2. Minkowski norms
For the necessary condition, use that the Hessian of $\parallel\cdot\parallel^{2}/2$ at 0, is well defined, symmetric and it agrees with $\parallel\cdot\parallel^{2}$, see [182, Prop. 4.1]. ∎
Then a natural regularity for $\parallel\cdot\parallel$ is smoothness (at least $C^{2}$) away from 0 and, then, so will be its indicatrix $S=\parallel\cdot\parallel^{-1}(1)$. In this case, putting $L:=\parallel\cdot\parallel^{2}$ the fundamental tensor of the norm is:
$g_{v}:=\frac{1}{2}\,\hbox{Hess}\,_{v}L\qquad\forall v\neq 0.$ (1)
@@
A Minkowski norm $L_{0}=\parallel\cdot\parallel_{0}^{2}$ is a non-symmetric norm (according to Prop. 2.1, now $\parallel\cdot\parallel_{0}$ positive homogeneous and $L_{0}$ 2-homogeneous) which is smooth outside 0 and has a strongly convex indicatrix $\Sigma_{0}=L_{0}^{-1}(0)$, that is, its fundamental tensor defined as in (1) is positive definite.
The following property of Minkowski norms will rely on our choice of convexity. For any non-symmetric norm $\parallel\cdot\parallel$ smooth away from 0, the Legendre or (non-linear) flat map is
$\begin{array}[]{ccc}\flat:V\rightarrow V^{*},&&v\mapsto v^{\flat}:=(dL)_{v}=g_{v}(v,\cdot).\end{array}$ (2)
@@
This map is clearly continuous, positive homogeneous and onto. It also extends the standard lowering index map of scalar products. However it may be non linear and (see [52, Sect. 1.2]):
@@ 2.1.3. Wind norms
Coming back to our original problem on propagation §2.1.1, we can wonder what happens if the 0 velocity is not included in the open domain $D$, that is, in the cases of either critical wind, when $0\in S=\partial D$ or strong wind, when $0\not\in\bar{D}:=D\cup\partial D$. In the latter case, a natural cone with vertex at 0 and tangent to $S$ emerges and $S$, up to the tangency points, is the union of two open connected subsets $S_{\hbox{\small{conv}}}$ and $S_{\hbox{\small{conc}}}$ (see Fig. 6).
In a natural way, $S_{\hbox{\small{conv}}}$ can be regarded as the indicatrix of a conic Minkowski norm defined only in the interior of the cone. It also gives a strict triangle inequality for vectors in the interior domain of the cone (see Def. 2.14 and Section 2 in [104]). Analogously, $S_{\hbox{\small{conc}}}$ gives a Lorentz norm and a similar reverse triangle inequality holds (consider a similar picture as in Fig. 4 with a piece of concave indicatrix), see [107, §3.1, Appendix B]. In both cases, we can define the square of the norm $L_{0}$ and, thus, the fundamental tensor $g_{v}$ (as in (1)) for $v\neq 0$ in the open conic region, by means of the positive homogeneous extension:
@@
In the Lorentz part, the reverse triangle inequality has a natural interpretation even in a classical setting. Namely, in our example, the Lorentz norm correspond to the case when the zeppelin engine is trying to move in the direction opposite to the wind (describing a straight line in the conic region), however, as it is drifted by the wind, the arrival time is maximized.
@@
Figure 6. Case of strong wind: emergence of convex and concave conic indicatrices in the interior of a cone.
@@
A wind Minkowski norm on $V$ is a smooth connected compact strongly convex hypersurface $\Sigma_{0}=\partial D_{0}$, where $D_{0}$ is the interior domain of $S_{0}$. Depending on whether $0\not\in\bar{D}_{0}$, $0\in\Sigma_{0}$ or $0\in D_{0}$ the wind is called strong, critical or weak, resp. (the latter being a Minkowski norm too).
@@ 2.1.4. Rheonomic emergence of cones
    Conic: $v\in\mathcal{C}_{0}\Rightarrow\lambda v\in\mathcal{C}_{0},\forall\lambda>0$.
@@
    Convex interior: $\mathcal{C}_{0}=\partial A_{0}$ in $V\setminus\{0\}$ with $A_{0}\subset V\setminus\{0\}$ an open, convex subset (as around Fig. 3), called the interior (or cone domain) of the cone.
@@
  • A vector hyperplane $H\subset V$ is called lightlike (resp. timelike, causal or spacelike) if $H$ is tangent to $\mathcal{C}_{0}$ (resp. contains a timelike vector, lies in one of the previous cases, lies in none of them).
  • A one-form $\omega:V\rightarrow\mathds{R}$ is called timelike (resp. causal) if $\omega(w)>0$ for all $w\in\bar{A}_{0}$ (resp. $\omega(w)>0$ for all $w\in A_{0}$). When $\omega$ is causal but non-timelike then it is called lightlike (and ker $\omega$ must be a lightlike hyperplane). $\bar{A}^{*}_{0},A^{*}_{0}$ and $\mathcal{C}^{*}_{0}$ will denote the set of causal, timelike and lightlike forms, resp.
@@
(1) The assumptions in the definition of cone can be widely optimized, see [107, §2.1].
@@
(3) Later, causality relations for $\mathcal{C}_{0}$ will behave as if v-v were “past directed”. As a consequence, the natural definition for a vector $v$ (in $V$ or $V\setminus\{0\}$) to be spacelike would be that neither $v$ nor v-v is causal.
(4) One can check that a vector hyperplane $H$ is spacelike if and only if for one (and then any) causal vector $v$ the affine hyperplane $v+H$ intersects transversely all the directions of $\mathcal{C}_{0}$ (then, yielding a strongly convex hypersurface in $H$ whenever $n>2$); moreover, any cone must admit spacelike hyperplanes [107, Lemma 2.5, Prop. 2.6].
@@ 2.1.5. Lorentz-Minkowski norms
In §2.1.3, the possibility to define a Lorentz norm associated with a concave indicatrix and a reverse triangle inequality was explained. In the particular case considered therein, the fundamental tensor $g_{v}$ had Lorentzian signature but it was defined only in the (timelike) directions in the interior of the cone (see (3) and Fig. 6). Indeed the Lorentz part of the norm could not be extended even continuously by 0 to the cone $\mathcal{C}_{0}$, because the indicatrix touched tangently the cone. Nevertheless, in the case of a Lorentzian scalar product the unit timelike vectors in one of the cones $\mathcal{C}_{0}$ do yield a Lorentz norm which can be extended by 0 to $\mathcal{C}_{0}$ and such that the corresponding fundamental tensor $g_{v}$ is well defined on $\mathcal{C}_{0}$ and non-degenrate (thus, with Lorentzian signature therein). Such properties are trivial in the case of a scalar product because this product coincides with $g_{v}$ whenever defined, but they serve as a guide for Lorentz norms with convenient regularity.
@@
Again, the hypotheses are not optimized, namely, (3) implies that $\mathcal{C}_{0}$ is a (strongly convex) cone even if one only assumes that $\mathcal{C}$ is the boundary of a domain invariant by positive homotheties (see [107, §3.1]). In any case, the indicatrix $\Sigma_{0}:=L^{-1}_{0}(1)$ will behave qualitatively as in Lorentz-Minkowski and it is asymptotic to the cone $\mathcal{C}_{0}$ (Fig. 8).
Remark 2.9.
As explained in §2.1.3, any Lorentz-Minkowski norm $L_{0}$ satisfies the strict reversed triangle inequalty
$\sqrt{L_{0}(v+w)}\geq\sqrt{L_{0}(v)}+\sqrt{L_{0}(w)},\qquad\forall v,w\in A_{0},$
with equality if and only $v,w$ are colinear (then necessarily pointing in the same orientation). Indeed, this property relied in the strict concaveness of the indicatrix. It is worth pointing out that this reverse triangle inequality permits to re-prove and extend some known inequalities (Aczél, Popoviciu, Bellman), see [129].
@@
Let $L_{0}:\bar{A}_{0}\rightarrow\mathds{R}$ be a Lorentz-Minkowski norm and let $\bar{A}^{*}_{0}$ be the set of all the causal forms (see Def. 2.6). Its Legendre (or flat) map is defined formally as in (2) adapting the domain and codomain, that is:
$\flat:\bar{A}_{0}\rightarrow\bar{A}^{*}_{0}\qquad(dL_{0})_{v}=v\rightarrow v^{\flat}:=g_{v}(v,\cdot).$
Reasoning as around (2), the indicatrix $\Sigma_{0}$ yields now a positive-homogeneous diffeomorphism between timelike vectors and forms which extend to lightlike ones, making the inverse $\sharp$ well defined. Notice, however, that for $\in\Sigma_{0}, vv^{\flat}$ can be identified with the affine hyperplane $v+T_{v}\Sigma_{0}$, while for $u\in\mathcal{C}_{0}$ the map $u\mapsto u+T_{u}\mathcal{C}_{0}$ does not yield and analogous identification (as all the radial affine hyperplanes $\lambda u+T_{\lambda u}\mathcal{C}_{0},\lambda>0,$ are equal).